# 37N35, 65F55, 90C22, 90C59, 93B52

### v1

@@ 3.4 Algorithm description
3Initialize $U_{0}=[\leavevmode\nobreak\ ]$, $X_{0}=[\leavevmode\nobreak\ ]$, $Y_{0}=[\leavevmode\nobreak\ ]$, $V_{0}=[\leavevmode\nobreak\ ]$ and $K_{0}=0$.
@@
14      Query the model $f$ at $x^{\operatorname{n}}(t_{j-1})$ and $u^{\operatorname{n}}_{j-1}(t_{j-1})$ to obtain $y^{\operatorname{n}}(t_{j-1})=f(x^{\operatorname{n}}(t_{j-1}),u^{\operatorname{n}}_{j}(t_{j-1}))$.
@@
holds for any data matrix $X$ and fixed feedback $K$; see [37]. However, the data generated by Algorithm 2 is typically well suited nevertheless for system identification. First, only $p$ additional iteration steps without subspace extension are needed to satisfy the rank conditions of the data matrices over the low dimensional subspace. These additional steps need to generate independent data, which can typically obtained using random input signals. Second, as for the construction of controllers based on linearization theory, data must be generated close to the steady state of interest to yield accurate results in linear system identification. Therefore, the use of stabilizing controllers in the data generation results in the desired behavior of trajectories to remain in the vicinity of the steady state.
The ICI method as shown in Algorithm 2 contains a numerical extension in Algorithm 2 that aims to improve the numerical stability of the method and allows for the constructed data triplets to satisfy the assumptions made in Theorem 2. This will be discussed in the following section.
@@ 3.5 Numerical stability and re-projections
4Initialize $U^{\operatorname{n}}=[\leavevmode\nobreak\ ]$, $X^{\operatorname{n}}=[\leavevmode\nobreak\ ]$, $Y^{\operatorname{n}}=[\leavevmode\nobreak\ ]$ and $k=1$.
@@ 27      Update the data matrices
By construction of the data triplet, we have that the row rank of $X^{\operatorname{n}}-\bar{x}$ is $r$. Also, for all linear systems, whose dynamics evolve in $\mathcal{V}$, we have that $n_{\min}\leq r$ holds for the intrinsic system dimension, and via the assumptions, it holds that $V^{\mkern-1.5mu\mathsf{T}}K=\widehat{K}$ is a stabilizing controller for all these systems. Therefore, we can apply Theorem 1 to the systems restricted to the subspace $\mathcal{V}$, which proofs that $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is informative for local stabilization of $f$ over $\mathcal{V}$ with only $T=r$ many data samples. ∎
Theorem 3 states that if a controller, which stabilizes the underlying system over $\mathcal{V}$, has been given to the re-projection method in Algorithm 3, the re-projected data triplet will be informative for stabilization over $\mathcal{V}$ already at step $r$ of the re-projection scheme. In other words, the number of data samples in the re-projected data set can be reduced from $T\geq r$ to $r$ while preserving their informativity. This is expected to increase the numerical stability as well as to reduce the computational costs of the ICI method. However, in particular in the presence of numerical errors, the application of more than $r$ re-projection steps may be desired to further numerically stabilize the low-dimensional controller inference. Therefore, in Algorithm 3 of Algorithm 3 additional data may be generated from random samples in $\mathcal{V}$.
The re-projection method from Algorithm 3 is incorporated into ICI (Algorithm 2) as a fall back if no stabilizing controller could be inferred from the current data triplet. This is inspired by the observation that in many cases when numerical instabilities cause the controller inference via linear matrix inequalities to fail computationally, the data quality improvements via re-projection can resolve this issue. Also, note that in this setup, re-projection is not applied in every iteration step, which keeps the number of overall model queries small. To further improve efficiency, we have added a testing procedure to Algorithm 3 that employs the theory of Theorem 3. If the number of requested data samples $T$ in Algorithm 3 is larger than the current subspace dimension $r$, we apply the low-dimensional controller inference method (Algorithm 1) after the $r$-th step of Algorithm 3 to test, if there exists a stabilizing controller $K$ over $\mathcal{V}_{j+1}$. In this case, the size of the data triplet is reduced from $T$ to $r$ for the next iteration of ICI.
@@ 4.1 Experimental setup and summary of results
Figure 1: In all experiments, the new iterative controller inference (ICI) method needs less system queries to generate data sets informative enough for stabilization of the underlying system, with improvement factors up to $25$. In the case of the power network example, we were not able to find a setup in which classical low-dimensional controller inference with unguided data sampling was able to stabilize the system.
We compare our proposed ICI method from Algorithm 2 to the non-adaptive low-dimensional controller inference (Algorithm 1; see [43]) using either a single trajectory of data samples or multiple trajectories, where the simulations have been restarted. In either case, random input signals drawn from a normal distribution with expectation $\mathbb{E}(u^{\operatorname{n}}(t))=\bar{u}$ have been used to excite the systems for the data generation. Due to numerical instabilities using the generated raw data for controller inference, also the non-adaptive approaches have been equipped with the re-projection scheme Algorithm 3.
@@ 4.2 Unstable heat flow
(b) iterative controller inference
(c) inferred from single trajectory
(d) inferred from multiple trajectories
Figure 3: Discrete-time unstable heat flow: The new ICI approach stabilizes the system after only $14$ system evaluations, which is around $1.6$ times less evaluations needed by either of the non-adaptive approaches. Low-order controller inference from a single or four trajectories using comparable numbers of system evaluations fail to stabilize the system.
[…]
The results for the continuous-time version of the system can be seen in Figure 2. The ICI method uses $28$ samples generated via the adaptive time integration routine ode45 leading overall to around 4 02440244\,0244 024 model evaluations. The final simulations of the uncontrolled and stabilized system in Figure 2 are computed using the ode15s solver to illustrate the independence of the computed controller from the chosen time discretization scheme in the continuous-time case.
@@
We then used the IMEX Euler scheme with the sampling time $\tau=0.1$ to discretize the linear system with the state dynamics implicitly in time and the inputs explicitly in time. The results for the three compared methods using similar numbers of model queries are shown in Figure 3, with ICI using $14$ queries, the controller inference from a single trajectory using $14$ model queries and the controller inference from four trajectories using $16$ queries. Only ICI is able to provide a stabilizing controller for the system with this small amount of evaluations, while the other two approaches fail. By increasing the number of model evaluations, the other methods provide stabilizing controllers after $22$ queries with a single trajectory and $24$ evaluations using four trajectories.
@@ 4.3 Brazilian interconnected power system
(b) iterative controller inference
Figure 5: Discrete-time power network: In the uncontrolled simulation, the output trajectories tend to infinity due to the instability of the power network. ICI provides a stabilizing controller after $20$ queries. The comparing low-dimensional controller inference setups did not provide a stabilizing controller within our limit of $300$ model queries.
[…]
The simulation with the controller constructed from ICI with $20$ model queries is shown in Figure 5. The low-dimensional controller inference based on a single or multiple trajectories using also $20$ model queries was not able to stabilize the system. These simulation results are not shown here but can be found in the accompanying code package [41]. In fact, we were not able to find any setup within our limit of $300$ model queries in which the non-adaptive low-dimensional controller inference was able to provide a stabilizing controller with either a single or multiple trajectories.
@@ 4.4 Tubular reactor
[…]
The continuous-time and discrete-time uncontrolled simulations and the results for the new ICI approach are shown in Figure 6. In the continuous-time case, for ICI we used the adaptive ode15s solver to generate data at $6$ discrete time points. The simulation of the resulting controller is shown in Figure 6(b). In the discrete-time case, ICI needed $8$ model queries to construct a stabilizing controller. The non-adaptive low-dimensional controller inference based on random inputs using the same sample number of model evaluations cannot stabilize the system. These results can be found in the accompanying code package [41]. However, increasing the number of allowed model queries to $12$ yields stabilizing controllers via a single or two trajectories used in the low-dimensional controller inference approach. This is a 1.5×1.5$\times$1.5 × increase in the number of model queries compared to ICI.
@@ 4.5 Crystal cluster growing
(b) iterative controller inference
(c) inferred from single trajectory
(d) inferred from multiple trajectories
Figure 8: Discrete-time crystal cluster: The three clusters quickly break apart and diverge from each other in the uncontrolled simulation and only the ICI approach yields a suitable controller using $16$ model evaluations. About $\times4$ more queries than in ICI are needed using an unguided single trajectory.
[…]
The simulations of the model can be seen in Figure 8. The outputs shown are the mean velocities of the particles of the clusters and any blow-up behavior implies that the clusters are drifting apart indefinitely. ICI provides a stabilizing controller after $16$ model queries. The non-adaptive low-dimensional controller inference using the same number of queries with a single or multiple trajectories from random inputs are not stabilizing as shown in Figures 8(c) and 8(d). Using a single trajectory only, $70$ model evaluations are needed to construct a stabilizing controller, which is around $\times4$ more queries than ICI needs. Allowing data from three trajectories reduces the number of model queries for the design of a stabilizing controller to $30$, which is still around 1.8×1.8$\times$1.8 × more than needed for ICI.
4.6 Laminar flow an obstacle
@@
(b) iterative controller inference
Figure 9: Laminar flow behind an obstacle: The upper plots show the velocity field in the last time step and the lower plots the mean flow velocities at the end of the channel. ICI can stabilizes the system using only $10$ system queries, while for non-adaptive controller inference 25×25$\times$25 × or 12×12$\times$12 × more model queries are needed to construct stabilizing controllers, depending on the use of only a single or six state trajectories.
[…]
The uncontrolled simulation and the results for the ICI-based feedback controller are shown in Figure 9 with snapshots of the velocity fields from the last time step and trajectories of the mean velocities at the end of the channel. ICI needed only $10$ model queries to provide a stabilizing controller for the system. The non-adaptive low-dimensional controller inference destabilized the system. Instead using either $250$ model queries in a single state trajectory or $120$ for multiple trajectories allowed also with this non-adaptive approach to construct stabilizing controllers. These numbers of model queries are $25$ and $12$ times larger than for ICI, respectively.
@@ 5 Conclusions
We introduced an adaptive sampling scheme for the construction of informative data sets for the task of constructing stabilizing feedback controllers. The approach is inspired by sufficient conditions for data informativity using stabilizing control signals as system inputs. Thereby, a sequence of stabilizing controllers is constructed over nested low-dimensional subspaces based on state data samples that are obtained by exciting the system using previous stabilizing controllers. The numerical experiments that up to an order of magnitude fewer samples are necessary than with unguided data collection. The method as presented in this work solely relies on forward evaluations of the underlying model. Future work includes combining the proposed adaptive sampling approach with adjoint information, which can further reduce the required number of data samples [42]. This is of particular interest in applications, where the instabilities affect high-dimensional subspaces such as in chaotic systems.

### v2

@@ 3.4 Algorithm description
3Initialize $U_{0}=[\penalty 10000\ ]$, $X_{0}=[\penalty 10000\ ]$, $Y_{0}=[\penalty 10000\ ]$, $V_{0}=[\penalty 10000\ ]$ and $K_{0}=0$.
@@
14  Query the model $f$ at $x^{\operatorname{n}}(t_{j-1})$ and $u^{\operatorname{n}}_{j-1}(t_{j-1})$ to obtain $y^{\operatorname{n}}(t_{j-1})=f(x^{\operatorname{n}}(t_{j-1}),u^{\operatorname{n}}_{j-1}(t_{j-1}))$.
@@
holds for any data matrix $X$ and fixed feedback $K$; see [37]. However, the data generated by Algorithm 2 is typically well suited nevertheless for system identification. First, only $p$ additional iteration steps without subspace extension are needed to satisfy the rank conditions of the data matrices over the low dimensional subspace. These additional steps need to generate independent data, which can typically be obtained using random input signals. Second, as for the construction of controllers based on linearization theory, data must be generated close to the steady state of interest to yield accurate results in linear system identification. Therefore, the use of stabilizing controllers in the data generation results in the desired behavior of trajectories to remain in the vicinity of the steady state.
The ICI method as shown in Algorithm 2 contains a numerical extension in Line 2 that aims to improve the numerical stability of the method and allows for the constructed data triplets to satisfy the assumptions made in Theorem 2. This will be discussed in the following section.
@@ 3.5 Numerical stability and re-projections
4Initialize $U^{\operatorname{n}}=[\penalty 10000\ ]$, $X^{\operatorname{n}}=[\penalty 10000\ ]$, $Y^{\operatorname{n}}=[\penalty 10000\ ]$ and $k=1$.
@@ 27  Update the data matrices
29  -
30 if $k+1\geq r$ and stabilizing $K$ can be inferred from $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ then
31    break
32   end if
33
@@
By construction of the data triplet, we have that the row rank of $X^{\operatorname{n}}-\bar{x}$ is $r$. Also, for all linear systems, whose dynamics evolve in $\mathcal{V}$, we have that $n_{\min}\leq r$ holds for the intrinsic system dimension. Thus, using the assumptions of this theorem, it holds that $KV=\widehat{K}$ is a stabilizing controller for all linear systems whose dynamics evolve in $\mathcal{V}$ and that are consistent with the given shifted data. Therefore, we can apply Theorem 1 to the systems restricted to the subspace $\mathcal{V}$, which proves that $(U^{\operatorname{n}},X^{\operatorname{n}},Y^{\operatorname{n}})$ is informative for local stabilization of $f$ over $\mathcal{V}$ with only $T=r$ many data samples. ∎
Theorem 3 states that if a controller, which stabilizes the underlying system over $\mathcal{V}$, has been given to the re-projection method in Algorithm 3, the re-projected data triplet will be informative for stabilization over $\mathcal{V}$ already at step $r$ of the re-projection scheme. In other words, the number of data samples in the re-projected data set can be reduced from $T\geq r$ to $r$ while preserving their informativity. This is expected to increase the numerical stability as well as to reduce the computational costs of the ICI method. However, in particular in the presence of numerical errors, the application of more than $r$ re-projection steps may be desired to further numerically stabilize the low-dimensional controller inference. Therefore, in Line 3 of Algorithm 3 additional data may be generated from random samples in $\mathcal{V}$.
The re-projection method from Algorithm 3 is incorporated into ICI (Algorithm 2) as a fall back if no stabilizing controller could be inferred from the current data triplet. This is inspired by the observation that in many cases when numerical instabilities cause the controller inference via linear matrix inequalities to fail computationally, the data quality improvements via re-projection can resolve this issue. Also, note that in this setup, re-projection is not applied in every iteration step, which keeps the number of overall model queries small. To further improve efficiency, we have added a testing procedure to Algorithm 3 in Line 3 that employs the theory of Theorem 3. If the number of requested data samples $T$ in Algorithm 3 is larger than the current subspace dimension $r$, we apply the low-dimensional controller inference method (Algorithm 1) after the $r$-th step of Algorithm 3 to test if there exists a stabilizing controller $K$ over $\mathcal{V}$ and to terminate the re-projection scheme early. In this case, the size of the data triplet is reduced from $T$ to $r$ for the next iteration of ICI.
@@ 4.1 Experimental setup and summary of results
Figure 1: In all experiments, the new iterative controller inference (ICI) method needs less system queries to generate data sets informative enough for stabilization of the underlying system, with improvement factors up to $25$. In the case of the power network example, we were not able to find a setup in which classical low-dimensional controller inference (ContInf) with unguided data sampling was able to stabilize the system.
We compare our proposed ICI method from Algorithm 2 to the non-adaptive low-dimensional controller inference (ContInf) from Algorithm 1 using either data samples from a single trajectory or multiple trajectories, where the simulations have been restarted. In either case, random input signals drawn from a normal distribution with expectation $\mathbb{E}(u^{\operatorname{n}}(t))=\bar{u}$ have been used to excite the systems for the data generation. Due to numerical instabilities using the generated raw data for controller inference, also the non-adaptive approaches have been equipped with the re-projection scheme Algorithm 3.
@@ 4.2 Unstable heat flow
The unstable heat flow example models the effects of overheating in machine tools. We consider here the HF2D5 model from [18], in which the heat equation is discretized in a two-dimensional rectangular domain. Disturbances are applied to the Laplace operator leading to instabilities. The example is described by a model of N=4 489N=4\,489 linear ODEs with $p=2$ inputs. The desired steady state is considered to be $(\bar{x},\bar{u})=(0,0)$ and the test simulations are performed with a constant input signal.
@@
The results for the continuous-time version of the system can be seen in Figure 2. The ICI method uses $28$ samples generated via the adaptive time integration routine ode45 leading overall to around 4 0244\,024 model evaluations. The final simulations of the uncontrolled and stabilized system in Figure 2 are computed using the ode15s solver to illustrate the independence of the computed controller from the chosen time discretization scheme in the continuous-time case.
@@
(b) ICI controller
(c) ContInf (single trajectory)
(d) ContInf (multiple trajectories)
Figure 3: Discrete-time unstable heat flow: The new ICI approach stabilizes the system after only $14$ system evaluations, which is around $1.6$ times less evaluations needed by either of the non-adaptive approaches. ContInf from a single or four trajectories using comparable numbers of system evaluations fail to stabilize the system.
@@
We then used the IMEX Euler scheme with the sampling time $\tau=0.1$ to discretize the linear system with the state dynamics implicitly in time and the inputs explicitly in time. The results for the three compared methods using similar numbers of model queries are shown in Figure 3, with ICI using $14$ queries, ContInf from a single trajectory using $14$ model queries and ContInf from four trajectories using $16$ queries. Only ICI is able to provide a stabilizing controller for the system with this small amount of evaluations, while the other two approaches fail. By increasing the number of model evaluations, the other methods provide stabilizing controllers after $22$ queries with a single trajectory and $24$ evaluations using four trajectories.
@@ 4.3 Brazilian interconnected power system
(b) ICI controller
Figure 5: Discrete-time power network: In the uncontrolled simulation, the output trajectories tend to infinity due to the instability of the power network. ICI provides a stabilizing controller after 2020 queries. The comparing low-dimensional controller inference setups did not provide a stabilizing controller within our limit of $300$ model queries.
This example is a linearized model of the Brazilian interconnected power network, which is unstable under heavy load conditions [20, 10]. We consider the model data bips98_606 from [7], which yields a system of linear differential-algebraic equations (DAEs) with N=7 135N=7\,135 and $p=4$. To simulate this DAE system, we discretized the model using the IMEX Euler scheme with the sampling time $\tau=0.1$.
The simulation with the controller constructed from ICI with $20$ model queries is shown in Figure 5. ContInf based on a single or multiple trajectories using also $20$ model queries was not able to stabilize the system. These simulation results are not shown here but can be found in the accompanying code package [41]. In fact, we were not able to find any setup within our limit of $300$ model queries in which the non-adaptive ContInf was able to provide a stabilizing controller with either a single or multiple trajectories.
@@ 4.4 Tubular reactor
This example describes an exothermic reaction inside a tubular reactor with one reactant [15, 46]. Due to the influence of the temperature on the reaction speed, the system enters an oscillating limit cycle behavior in the case of any disturbances. Following the one-dimensional spatial discretization in [16], we have a system of N=3 998N=3\,998 nonlinear ODEs with $p=2$ inputs that allow to control the reactant concentration at the beginning of the tube and the temperature along the tube. We consider the continuous-time as well as a discrete-time version of this example. The discrete-time model has been obtained using the IMEX Euler scheme with sampling time $\tau=0.01$, where the linear components are evaluated implicitly in time and the nonlinear components and controls explicitly in time.
The continuous-time and discrete-time uncontrolled simulations and the results for the new ICI approach are shown in Figure 6. In the continuous-time case, for ICI we used the adaptive ode15s solver to generate data at $6$ discrete time points. The simulation of the resulting controller is shown in Figure 6(b). In the discrete-time case, ICI needed $8$ model queries to construct a stabilizing controller. The non-adaptive ContInf based on random inputs using the same sample number of model evaluations cannot stabilize the system. These results can be found in the accompanying code package [41]. However, increasing the number of allowed model queries to $12$ yields stabilizing controllers via a single or two trajectories used in the ContInf approach. This is a $1.5\times$ increase in the number of model queries compared to ICI.
@@ 4.5 Crystal cluster growing
(b) ICI controller
(c) ContInf (single trajectory)
(d) ContInf (multiple trajectories)
Figure 8: Discrete-time crystal cluster: The three clusters quickly break apart and diverge from each other in the uncontrolled simulation and only the ICI approach yields a suitable controller using $16$ model evaluations. About $4.4\times$ more queries than in ICI are needed by ContInf using unguided single trajectory data.
The Toda lattice model is used to describe one-dimensional crystals [35, 40] using classical physics laws from solid-state mechanics about mass-spring-damper systems. Here, we consider the case of a crystalization process in which three major crystal clusters have been formed already. Stabilization is needed to hold these clusters in place to allow for the full crystal to form; see [44] for further details of the setup including the describing equations. In our experiments, we are using 2 0002\,000 particles leading to a system of N=4  [^0] N=4\,000 nonlinear ODEs with $p=3$ inputs to change the displacement of the three clusters. For the discretization of the model, we use the IMEX Euler scheme with sampling time $\tau=0.1$, where the linear parts are handled implicitly and the nonlinear parts and controls explicitly.
The simulations of the model can be seen in Figure 8. The outputs shown are the mean velocities of the particles of the clusters and any blow-up behavior implies that the clusters are drifting apart indefinitely. ICI provides a stabilizing controller after $16$ model queries. The non-adaptive ContInf using the same number of queries with a single or multiple trajectories from random inputs are not stabilizing as shown in Figures 8(c) and 8(d). Using a single trajectory only, $70$ model evaluations are needed to construct a stabilizing controller, which is around $4.4\times$ more queries than ICI needs. Allowing data from three trajectories reduces the number of model queries for the design of a stabilizing controller to $30$, which is still around $1.8\times$ more than needed for ICI.
4.6 Laminar flow behind an obstacle
@@
(b) ICI controller
Figure 9: Laminar flow behind an obstacle: The upper plots show the velocity field in the last time step and the lower plots the mean flow velocities at the end of the channel. ICI can stabilizes the system using only $10$ system queries, while for non-adaptive ContInf $25\times$ or $12\times$ more model queries are needed to construct stabilizing controllers, depending on the use of only a single or six state trajectories.
As final example, we consider the laminar flow behind an obstacle, which is described by the incompressible Navier-Stokes equations [2]. At medium Reynolds numbers—here $\operatorname{Re}=90$—the non-zero steady state tends to unstable oscillatory vortex shedding. We consider the example matrices given in [2], providing a spatially discretized model with N=6 618N=6\,618 nonlinear DAEs and $p=6$ inputs that allow the change of the flow velocities in a small rectangle behind the obstacle. For the time simulation, the system is discretized in time via the IMEX Euler scheme with sampling time $\tau=0.0025$ and where the linear parts are handled implicitly and the nonlinear components and controls explicitly.
The uncontrolled simulation and the results for the ICI-based feedback controller are shown in Figure 9 with snapshots of the velocity fields from the last time step and trajectories of the mean velocities at the end of the channel. ICI needed only $10$ model queries to provide a stabilizing controller for the system. In contrast, to stabilize the system with the non-adaptive ContInf approach, either $250$ model queries in a single state trajectory or $120$ queries for multiple trajectories are needed to construct stabilizing controllers. These numbers of model queries are $25$ and $12$ times larger than for ICI, respectively.
@@ 5 Conclusions
We introduced an adaptive sampling scheme for the construction of informative data sets for the task of constructing stabilizing feedback controllers. The approach is inspired by sufficient conditions for data informativity using stabilizing control signals as system inputs. Thereby, a sequence of stabilizing controllers is constructed over nested low-dimensional subspaces based on state data samples that are obtained by exciting the system using previous stabilizing controllers. The numerical experiments demonstrate that up to one order of magnitude fewer samples are needed by the proposed approach compared to unguided data collection. The method as presented in this work solely relies on forward evaluations of the underlying model. Future work includes combining the proposed adaptive sampling approach with adjoint information, which can further reduce the required number of data samples [42]. This is of particular interest in applications where the instabilities affect high-dimensional subspaces such as in chaotic systems.